Michelangelo Inc., a software development firm, has stock outstanding as follows: 30,000 shares of cumulative 1%, preferred stock of $20 par, and 38,000 shares of $75 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $2,400; second year, $3,300; third year, $47,260; fourth year, $124,940. Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".
Answer #1

Annual preferred dividend = 30000*20*1% = 6000

1st year 2nd year 3rd year 4th year
Preferred dividend (per share) 2400/30000 = 0.08 3300/30000 = 0.11 12300/30000 = 0.41 6000/30000 = 0.20
Common dividend (per share) 0 0 34960/38000 = 0.92 118940/38000 = 3.13